# Efficient simulation of the SABR model

### v1

@@ Appendix B Sampling Islah’s approximation via a CEV distribution
Though Islah’s approximation is less than desirable, we show that Islah’s distribution of $F_{t+h}$ can be expressed as a power of a CEV random variable. Therefore, the CEV simulation algorithm in Section 4.2 can be applied to expedite the sampling.
Let us define
@@
As a result, the complementary CDF of Islah (2009)’s approximation is equivalently expressed by
$\mathrm{Prob}(F_{t+h}\geq y)=\mathrm{Prob}\left(\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\geq y^{\prime}\right)=P_{\chi^{2}}\left(\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}};\frac{1}{\beta_{\ast}^{\prime}},\frac{\left(y^{\prime}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}}\right),$
@@
In other words, a power of $F_{t+h}$ instead of $F_{t+h}$ itself, follows a CEV distribution under Islah’s approximation. Therefore, $F_{t+h}$ can be sampled as

### v2

@@ Appendix B Sampling Islah’s approximation via a CEV distribution
Though Islah’s approximation of $F_{t+h}$ is less than desirable since it fails the martingale property, nevertheless, we show that Islah’s approximation can be expressed as a power function of a CEV random variable. For theoretical interest, we show that the CEV simulation algorithm in Section 4.2 can be applied to expedite the sampling of $F_{t+h}$ based on Islah’a approximation.
We define
@@
As a result, the complementary CDF of Islah’s approximation is equivalently expressed by
$\mathrm{Prob}(F_{t+h}\geq y)=\mathrm{Prob}\left(\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\geq y^{\prime}\right)=P_{\chi^{2}}\left(\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}};\frac{1}{\beta_{\ast}^{\prime}},\frac{\left(y^{\prime}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\right),$
@@
In other words, instead of $F_{t+h}$ itself, a power function of $F_{t+h}$ follows a CEV distribution under Islah’s approximation. Therefore, $F_{t+h}$ can be sampled as